OpenStudy (anonymous):

Alex is painting an 8-foot long mural on the wall of the community center. If he has finished 1/4 of the mural, how many feet of the mural has he painted?

OpenStudy (anonymous):

PLEASE HELP!!

hero (hero):

He painted 1/4 *8 = 8/4 = 2 feet

OpenStudy (anonymous):

thnx :)

hero (hero):

You can also just take 8 and just divide it by 4 8/4 = 2

hero (hero):

If you're looking for 1/4 of a number, just take the number and divide by 4: 1/4 of n = n/4